What are the answers to these questions, and how do I answer them?

Below is are balance and income numbers for three companies. Answer the questions below. ​

1

2

3

Year 1

Year 2

Year 1

Year 2

Year 1

Year 2

Revenue

5674

5054

176896

124959

179083

186506

COGS

3442

3066

144924

102374

144249

150228

Gross Profit

2232

1988

31972

22585

34834

36277

EBITDA

1470

658

7962

2712

15990

20527

Taxes

510

218

6522

325

4309

11615

Net Income

960

440

1440

2387

11681

8912

Total Assets

5255

4681

275936

194921

244587

254725

Equity

3084

3235

13442

40922

89999

134333

Current Assets

3145

1550

62923

23413

91387

144135

Current Liabilities

873

289

53604

6438

85373

48307

Inventory

221

513

10271

2812

13799

81558

Long-Term Debt

1298

1156

208890

147560

69215

72084

  • ● The balance sheet items may not balance perfectly (they may be out ~$1) due to rounding error.

  • ● USE EXCEL to calculate the solutions and do the math – calculators may throw off your rounding

  • ● Submit percentages as the percentage followed by two decimals. A number that appears as .04567 in excel should be

    submitted as 4.57%.

  • ● Responses will be marked correct if they are within 5% of the answer I calculated in Excel

    (Questions on next page)

  1. Which company has the best liquidity (highest current ratio) in Year 2?

  2. Which company does the best job generating profits for shareholders in Year 2? (ROE)

  3. Which company does the best job generating profits with its total investment in Year 2? (ROA)

  4. Which company is the most leveraged (high long-term debt to equity) in Year 2?

  5. Which company has the biggest current threat of being illiquid (lowest current ratio) in Year 2?

  6. Which company saw the biggest increase in its net profit margin from Year 1 to Year 2 (or smallest decrease if NI all

    negative)?

  7. What is company 2's ROA in Year 2?

  8. What is company 1's ROA in Year 1?

  9. What is company 3's ROA in Year 2?

  10. Which company has the highest gross margin in Year 2?

  11. What is company 1's quick ratio in Year 1?

  12. What is company 2's current ratio in Year 2?

  13. What is company 3's net margin in Year 2?

  14. What is company 2's inventory turnover in Year 2?

  15. What is company 3's long-term debt to assets ratio in Year 1?
Answer #1

Current Ratio :

The Current Ratio compares all of a company's current assets to its current liabilities. Current Assets means cash and other assets which will turn into cash in a year or less. Current liabilities are those liabilities that will be paid in a year or less.

Formula: Current Ratio = Current Assets / Current Liabilities

Calculation of Companies having the Highest Current Ratio in Year - 2:

For Company -1: Current Ratio = 1550/289 = 5.36

For Company -2: Current Ratio = 23,413/6438 = 3.64

For Company -3: Current Ratio = 144,135/48,307 = 2.98

From the above calculation it is clear that Company - 1 is having highest liquidity ratio as it is having $ 5.63 as current asset for every $1 of current liability.

Return on Equity (ROE):

Return on Equity measures how effectively management is using a company's assets to create profits. It is calculated by diving Net Income by Shareholders' Equity.

Formula: Return on Equity = Net Income/Average Shareholders' Equity  

where, Average Shareholders' Equity = (Equity at the start of the Year+Equity at the end of the Year)/2.

Calculation of Companies having the Highest Return on Equity in Year - 2:

For Company -1: Return on Equity = 440/((3,084+3,235)/2) = 13.93%

For Company -2: Return on Equity = 2,387/((13,442+40,922)/2) = 8.78%

For Company -3: Return on Equity = 8,912/((89,999+134,333)/2) = 7.95%

From the above calculation it is clear that Company - 1 is having highest Return on Equity of 13.93%.

Return on Assets (ROA):

Return on Assets measures how effectively a company is utilizing its assets by calculating the percentage of profit a company earns in relation to its total assets.

Formula: Return on Assets = Net Income / Average Total Assets

where, Average Total Assets = (Total Assets at the start of the Year+Total Assets at the end of the Year)/2

Calculation of Companies having the Highest Return on Assets in Year - 2:

For Company -1: Return on Assets = 440/((5,255+4,681)/2) = 8.86%

For Company -2: Return on Assets = 2,387/((275,936+194,921)/2) = 1.01%

For Company -3: Return on Assets = 8,912/((244,587+254,725)/2) = 3.57%

From the above calculation it is clear that Company - 1 is having highest Return on Equity of 8.86%.

Debt to Equity Ratio:

Debt to Equity Ratio compares the company's total liabilities to its shareholder's equity and can be used to evaluate how much leverage a company is using. It is calculated by dividing the company's' total liabilities by its shareholder equity.

Formula: Debt to Equity Ratio = Total Liabilities / Total Shareholders' Equity

Calculation of Companies having the Highest Leverage i.e. Debt to Equity Ratio in Year - 2:

For Company -1: Debt to Equity Ratio = (289+1,156)/3235 = 0.45

For Company -2: Debt to Equity Ratio = (6,438+147,560)/40,922 = 3.76

For Company -3: Debt to Equity Ratio = (48,307+72,084)/134,333 = 0.90

From the above calculation, it is clear that Company - 2 is having highest leverage i.e. Debt to Equity Ratio of 3.76 which means that for every $ 1 of equity there is debt of $ 3.76.

Note: As per Chegg Policy we are allowed to answer maximum 4 sub-parts of the question.